oregonduck (oregonduck):

@ali2x2 Fiona has $18 to spend. She spent $4.25, including tax, to buy a notebook. She needs to save $9.75, but she wants to buy a snack. If crackers cost $0.50 per package including tax, what inequality would show the maximum number of packages that Fiona can buy? Solve your inequality by showing your work and steps.

OpenStudy (ali2x2):

starts with 18 spends 4.25 18 - 4.25 spends .50 on each cracker 18 - 4.25 - 0.50*(#crackers) all that should be greater than or equal to 9.75, the amount she needs to keep

OpenStudy (ali2x2):

18 - 4.25 - 0.50*(#crackers) is greater than or equal to 9.75 solve for (#crackers)

OpenStudy (ali2x2):

or, i could just tell you

oregonduck (oregonduck):

no it is fine so what do I put?

OpenStudy (ali2x2):

starts with 18 spends 4.25 left with 13.75 spends 0.50 per crackers ---0.50*(#crackers) wants at least 9.75 left over 13.75 - 0.50*(#crackers) >= 9.75

OpenStudy (ali2x2):

\[13.75−9.75≥0.5∗(crackers)\]

OpenStudy (ali2x2):

13.75 - 0.50*(#crackers) >= 9.75 add 0.50*(crackers) to each side 13.75 >= 9.75 + 0.5*(crackers) subtract 9.75 from both sides 4 >= 0.5*(crackers) divide by 0.5 on both sides 8 >= (crackers) eight is greater than or equal to the amount of crackers